# NVILA: Efficient Frontier Visual Language Models

### v1

@@ Abstract
Abstract:   Visual language models (VLMs) have made significant advances in accuracy in recent years. However, their efficiency has received much less attention. This paper introduces NVILA, a family of open VLMs designed to optimize both efficiency and accuracy. Building on top of VILA, we improve its model architecture by first scaling up the spatial and temporal resolutions, and then compressing visual tokens. This “scale-then-compress” approach enables NVILA to efficiently process high-resolution images and long videos. We also conduct a systematic investigation to enhance the efficiency of NVILA throughout its entire lifecycle, from training and fine-tuning to deployment. NVILA matches or surpasses the accuracy of many leading open and proprietary VLMs across a wide range of image and video benchmarks. At the same time, it reduces training costs by 4.5$\times$, fine-tuning memory usage by 3.4$\times$, pre-filling latency by 1.6-2.2$\times$, and decoding latency by 1.2-2.8$\times$. We will soon make our code and models available to facilitate reproducibility.
@@
Figure 1: Overview of NVILA: Efficient Frontier VLM. NVILA pioneers the development of highly efficient designs that span from the training phase to the deployment stage. Fig. (a,b,c): NVILA’s innovative techniques lead to a remarkable 4.5x acceleration in training speed, a significant 3.4x reduction in memory usage and 100x saving in parameters during fine-tuning. During inference, NVILA achieves a 1.8x improvement in Time to First Token and a 1.2x boost in throughput; Fig. (d): The efficient designs do no compromise the accuracy, instead, NVILA sustains competitive performance with existing SOTA models on both video and image benchmarks.
Figure 2: NVILA’s qualitative examples.
Figure 3: NVILA’s qualitative examples.
@@ 1 Introduction
Visual language models (VLMs) have shown remarkable abilities in processing and integrating both visual and textual information, enabling advanced vision-language interactions and dialogues. In recent years, the research community has made tremendous progress in enhancing their accuracy [1, 2, 3, 4, 5] and broadening their applications across diverse domains, including robotics [6, 7], autonomous driving [8], and medical imaging [9]. However, there has been much less focus on improving their efficiency.
@@
Thanks to these innovations, NVILA is both efficient and accurate. It reduces training costs by 4.5$\times$, fine-tuning memory usage by 3.4$\times$, pre-filling latency by 1.6–2.2$\times$, and decoding latency by 1.2–2.8$\times$. It also matches or surpasses the accuracy of many leading open VLMs [3, 4, 2] and proprietary VLMs [10, 11] across a wide range of image and video benchmarks. Furthermore, NVILA enables new capabilities including temporal localization, robotic navigation, and medical imaging. We will release our code and models soon to support full reproducibility. We hope our work will inspire further research on efficient VLMs.
@@ 2.1 Efficient Model Architecture
In this paper, we advocate for the “scale-then-compress” paradigm: first, we scale up the spatial/temporal resolutions to improve accuracy, and then we compress the visual tokens to improve efficiency. Scaling resolutions up improves the performance ceiling, but doing so alone will significantly increase the computational cost. For example, doubling the resolution will double the number of visual tokens, which will increase both training and inference costs by more than 2$\times$, as self-attention scales quadratically with the number of tokens. We can then cut this cost down by compressing spatial/temporal tokens. Compressed visual tokens have a higher information density, allowing us to preserve or even improve spatial and temporal details with fewer total tokens.
Table 1: Spatial “scale-then-compress”. Increasing the spatial resolution with Dynamic-S2 can greatly improve the model’s accuracy, particularly on text-heavy benchmarks. Compressing the visual tokens with spatial pooling can effectively reduce both the number of tiles and tokens per tile, with moderate accuracy loss. This loss can be further reduced by adding an additional visual encoder pre-training (VEP) stage. In this and following tables, “Image-10” refers to the average validation scores from the 10 benchmarks listed in Table 8.
Spatial Pooling #Tokens/Tile #Tiles/Image AI2D DocVQA TextVQA Image-10
@@ 2.2 Efficient Training
Figure 5: DeltaLoss visualizations in NVILA training: Left, Middle, and Right sections show examples that are too easy, distracting, and helpful for training, respectively.
@@ 2.2.1 Dataset Pruning
Table 3: Dataset pruning. DeltaLoss consistently rivals other data selection methods and shows negligeble performance drop when pruning 50% of data.
Method Image-10 MMMU DocVQA TextVQA
@@ 2.2.2 FP8 Training
Table 4: FP8 training. FP8 accelerates the training of NVILA while maintaining the accuracy, especially when gradient checkpointing (GC) is not enabled. In this table, the throughput results are obtained with the maximum achievable batch size (BS) on 64 H100 GPUs. VideoMME results come from an 8-frame setting and with subtitle information.
@@
FP16 and BF16 [28] have become standard precisions for model training, since they offer acceleration without accuracy loss, supported natively by NVIDIA GPUs. With the advent of the NVIDIA Hopper and Blackwell architectures, new GPUs (such as H100 and B200) now provide native support for FP8, which has emerged as a promising precision due to its potential for larger computational and memory efficiency.
@@
Table 5: Fine-tuning. PEFT Study for Efficient VLM Finetuning. (BS=1, #image=1, VE image resolution=448, patch size=14, LLM context length=300, gradient checkpointing disabled) Training time is profiled on NVIDIA A100-80G for one forward+backward pass.
ViT LLM Accuracy (Avg of 5) Memory (GB) Throughput (iter/s)
LoRA LoRA 71.8 20.1 3.4
LN LoRA 71.4 19.2 4.5
FT LoRA 70.1 21.9 4.2
LoRA QLoRA 70.8 11.1 2.6
LN QLoRA 70.9 10.2 3.1
FT FT 77.7 63.5 6.1
Table 6: Quantization. The impact of quantization on model’s accuracy. While W4A16 quantization on LLM backbone may introduce small accuracy drop, W8A8 quantization on ViT is nearly lossless.
AI2D MMMU VideoMME TTFT (s)
NVILA-8B 91.0 50.7 63.9 0.90
+ W4A16 LLM 90.9 49.2 62.0 0.77
+ W8A8 ViT 90.9 49.3 62.1 0.65
@@ 2.3 Efficient Fine-Tuning
When fine-tuning the vision encoder (ViT) and language model (LLM) together using PEFT methods, we observe that the learning rate should be set differently for VE and LLM: the learning rate for the ViT part will be 5-50$\times$ smaller than that for the LLM part. On the other hand, we also observe that fine-tuning the vision encoder with Layernorm can achieve comparable performance as LoRA (Table. 5) while being more computational-efficient: it can reduce the training time by 25% compared to applying LoRA for the vision encoder. With the curated configuration setup, NVILA can be quickly fine-tuned to various downstream tasks under 24 GB memory with on-par performance.
@@ 2.4 Efficient Deployment
We develop a specialized inference engine with quantization techniques to efficiently deploy NVILA. The inference process is divided into two phases: prefilling and decoding In the compute-bounded prefilling stage,, we first apply token compression techniques (Section 2.1) to reduce the inference workload for LLM backbone, after which the vision tower becomes the primary bottleneck, accounting for over 90% of the prefilling latency. To tackle this, we implement W8A8 quantization for the vision tower to reduce NVILA’s Time-To-First-Token (TTFT) in this compute-bounded stage. For the memory-bounded decoding stage, we follow AWQ [32] for W4A16 quantization of the LLM backbone to accelerate. We further optimize the original AWQ implementation by introducing FP16 accumulation to the W4A16 GEMM kernels, resulting to a total 1.7$\times$ kernel speedup without compromising accuracy. We attach a detailed comparison in Figure. 6
@@ 3.2 Accuracy Results
Table 8: Image benchmark results. We mark the best performance bold and the second-best underlined.
@@
test test test test testmini val/test/pro image val testdev
GPT-4o 94.2 85.7 92.8 79.2 63.8 69.1/64.7/51.9 75.4 76.2 77.4 78.7
Claude 3.5 Sonnet 94.7 90.8 85.2 74.3 67.7 68.3/63.7/51.5 60.1 74.1 70.7
Gemini 1.5 Pro 94.4 87.2 93.1 81.0 63.9 62.2/57.6/43.5 70.4 78.7 80.2
LLaVA-1.5 7B 55.5 17.8 28.1 25.8 25.6 35.7/–/– 54.8 66.1 58.2 78.5
VILA-1.5 8B 76.6 52.7 40.6 25.9 36.7 38.6/32.7/– 52.7 73.8 68.5 83.0
Cambrian-1 8B 73.0 73.3 77.8 41.6 49.0 42.7/–/– 64.2 74.7 71.7 81.2
LLaVA-OV 7B 81.4 80.0 87.5 68.8 63.2 48.8/42.8/24.1 66.3 75.4 78.3 84.0
Llama 3.2 11B 91.9 83.4 88.4 51.5 50.7/–/– 75.2
InternVL2 8B 83.8 83.3 91.6 74.8 58.3 51.2/42.6/29.0 64.2 76.2 77.4 76.7
Qwen2-VL 7B 83.0 83.0 94.5 76.5 58.2 54.1/46.6/30.5 70.1 84.3 82.9
NVILA-Lite 8B 91.0 84.8 91.7 67.9 64.5 50.7/45.7/26.5 65.6 76.3 78.1 85.0
NVILA 8B 92.3 86.1 93.7 70.7 65.4 49.9/44.4/27.8 68.6 76.5 80.1 85.4
LLaVA-1.5 13B 61.1 18.2 30.3 29.4 27.7 37.0/–/– 55.3 68.2 61.3 80.0
VILA-1.5 13B 79.9 59.5 58.6 30.4 42.7 37.9/33.6/– 57.5 72.6 65.0 82.8
Cambrian-1 13B 73.6 73.8 76.8 48.0 40.0/–/– 63.0 74.4 72.8
Pixtral 12B 79.0 81.8 90.7 50.8 58.0 52.5/–/– 65.4 75.7 80.2
NVILA-Lite 15B 92.0 81.8 90.6 69.3 61.7 58.7/51.8/33.7 67.1 75.6 77.3 83.7
NVILA 15B 94.1 86.9 94.0 73.5 66.1 56.7/51.8/33.8 69.5 76.6 80.0 84.8
VILA-1.5 40B 88.9 67.8 58.6 38.4 49.3 51.9/46.9/25.0 60.8 69.1 73.6 84.3
Cambrian-1 34B 79.7 75.6 75.5 46.0 53.2 49.7/–/– 67.8 75.3 76.7 83.8
LLaVA-NeXT 34B 46.5 48.1/44.5/22.9 75.9 69.5 83.7
InternVL2 40B 87.1 86.2 93.9 78.7 63.7 55.2/47.4/34.2 71.8 78.2 83.0
LLaVA-OV 72B 85.6 83.7 91.3 74.9 67.5 56.8/52.3/31.0 71.9 75.4 80.5 85.2
Llama 3.2 90B 92.3 85.5 90.1 57.3 60.3/–/39.5
NVLM-D-1.0 78B 94.2 86.0 92.6 65.2 59.7/54.6/– 69.7 82.1 85.4
@@ 3.2.1 Image Benchmarks
Our NVILA performs comparably to top open-source models in each size category, including Qwen2-VL [3], InternVL [4], and Pixtral. For general visual question answering tasks (ChartQA, DocVQA, InfoVQA, TextVQA, VQAv2, Seed), NVILA-8B and NVILA-15B achieve competitive or even better results compared to proprietary models (GPT-4o, Gemini). On science related benchmarks (AI2D), NVILA-8B achieves state-of-the-art performance among open-source models. When scaling to 15B, NVILA demonstrates competitive performance with proprietary models.
Further, on reasoning and knowledge benchmarks such as MMMU, RealworldQA, and MathVista, scores improves more when the model size increase. For benchmarks that require OCR capability such as TextVQA, AI2D, ChartQA, DocVQA, InfoVQA, 8B model can also do a great job. We also show a few qualitative examples in Figure. 2-3 to demonstrate the OCR, reasoning and multi-image capability of NVILA model.
@@ 3.2.2 Video Benchmarks
We evaluate our models on a range of video understanding benchmarks, spanning short videos of a few seconds to longer videos up to an hour in duration. Table 9 presents the performance of NVILA compared to baseline models. NVILA features long-context capability and can process up to 256 frames. With the scale-then-compress design, NVILA-8B achieves impressive results, setting new state-of-the-art performance across all benchmarks. Notably NVILA reaches performance levels comparable to GPT-4o mini with only 8B parameters and outperforms many larger models.
Table 9: Video benchmark results.
@@
LLaVA-OneVision 7B 32 56.6 56.5 64.7 56.7 79.4 58.2 61.5
Oryx-1.5 7B 128 56.3 67.5 67.6 81.8 58.8 64.2

### v2

@@ Abstract
Abstract:   Visual language models (VLMs) have made significant advances in accuracy in recent years. However, their efficiency has received much less attention. This paper introduces NVILA, a family of open VLMs designed to optimize both efficiency and accuracy. Building on top of VILA, we improve its model architecture by first scaling up the spatial and temporal resolutions, and then compressing visual tokens. This “scale-then-compress” approach enables NVILA to efficiently process high-resolution images and long videos. We also conduct a systematic investigation to enhance the efficiency of NVILA throughout its entire lifecycle, from training to deployment. NVILA matches or surpasses the accuracy of many leading open and proprietary VLMs across a wide range of image and video benchmarks. At the same time, it reduces training costs by 1.9-5.1$\times$, prefilling latency by 1.6-2.2$\times$, and decoding latency by 1.2-2.8$\times$. We make our code and models available to facilitate reproducibility.
@@
Figure 1: NVILA – Efficient Frontier VLMs. (a) NVILA trains image and video models 5.1$\times$ and 1.9$\times$ faster, respectively, than LLaVA-OneVision (OV), which is the only baseline model with publicly disclosed training costs. (b) Against Qwen2-VL, NVILA achieves a 1.6–2.2$\times$ measured speedup in the pre-filling stage and a 1.2–2.8$\times$ speedup during the decoding stage. (c) NVILA’s efficiency is achieved without compromising accuracy; in fact, it delivers comparable or even superior accuracy across image and video benchmarks. All models in this table have 8B parameters. Training time in (a) is measured using NVIDIA H100 GPUs, while inference speed in (b) is measured using a single NVIDIA GeForce RTX 4090 GPU. Accuracy numbers in (c) are normalized relative to the highest score for each benchmark.
Figure 2: Qualitative examples.
@@ 1 Introduction
Visual language models (VLMs) have shown remarkable abilities in processing and integrating both visual and textual information, enabling advanced vision-language interactions and dialogues. In recent years, the research community has made tremendous progress in enhancing their accuracy [1, 2, 3, 4, 5] and broadening their applications across diverse domains, including robotics [6, 7, 8], autonomous driving [9], and medical applications [10, 11]. However, there has been much less focus on improving their efficiency.
@@
Thanks to these innovations, NVILA is efficient and accurate. It reduces training costs by 1.9-5.1$\times$, prefilling latency by 1.6–2.2$\times$, and decoding latency by 1.2–2.8$\times$. It also matches or surpasses the accuracy of leading open VLMs [5, 3, 2] and proprietary VLMs [12, 13] across a wide range of image and video benchmarks. Furthermore, NVILA enables new capabilities including temporal localization, robotic navigation, and medical imaging. We release our code and models to support full reproducibility. We hope our work will inspire further research on efficient VLMs.
@@ 2.1 Efficient Model Architecture
In this paper, we advocate for the “scale-then-compress” paradigm, where we first scale up the spatial/temporal resolutions to improve accuracy, and we then compress the visual tokens to improve efficiency. Scaling resolutions up improves the performance ceiling, but doing so alone will significantly increase the computational cost. For example, doubling the resolution will double the number of visual tokens, which will increase both training and inference costs by more than 2$\times$, as self-attention scales quadratically with the number of tokens. We can then cut this cost down by compressing spatial/temporal tokens. Compressed visual tokens have a higher information density, allowing us to preserve or even improve spatial and temporal details with fewer total tokens.
Table 1: Spatial “scale-then-compress”. Increasing the spatial resolution with Dynamic-S2 can greatly improve the model’s accuracy, particularly on text-heavy benchmarks. Compressing the visual tokens with spatial pooling can effectively reduce both the number of tiles and tokens per tile, with moderate accuracy loss. This loss can be further reduced by adding an additional visual encoder pre-training (VEP) stage. In this and following tables, “IM-10” refers to the average validation scores from the 10 benchmarks listed in Table 8.
Spatial Pooling #Tokens/Tile #Tiles/Image AI2D DocVQA TextVQA IM-10
@@ 2.2 Efficient Training
Figure 4: Dataset pruning. DeltaLoss visualizations in NVILA training: Left, Middle, and Right sections show examples that are too easy, distracting, and helpful for training, respectively.
@@ 2.2.1 Dataset Pruning
Table 3: Dataset pruning on NVILA Recipe. DeltaLoss consistently rivals other data selection methods and shows negligible performance drop when pruning 50% of data.
Method IM-10 MMMU DocVQA TextVQA
@@
We examine the impact of data pruning on newly added datasets. We incorporate varying percentages of pixmo data [30] into the NVILA training set. In Table LABEL:tab:dataset_prune_pixmo, we observe that directly combining pixmo data with the NVILA training set decreases performance on DocVQA and TextVQA benchmarks, while only improving MMMU scores. This suggests that aggressively increasing the training set size may actually hurt performance. By applying deltaloss to prune the training data—filtering out examples that are either too easy or too hard—we find that models trained with the pruned molmo dataset show general improvements in experimental results.
@@ 2.2.2 FP8 Training
Table 4: FP8 training. FP8 accelerates the training of NVILA while maintaining the accuracy, especially when gradient checkpointing (GC) is not enabled. In this table, the throughput results are obtained with the maximum achievable batch size (BS) on 64 H100 GPUs. Video-MME results come from an 8-frame setting and with subtitle information.
@@
FP16 [31] and BF16 [32] are standard precisions for model training, since they offer acceleration without accuracy loss, supported natively by NVIDIA GPUs. With the advent of the NVIDIA Hopper and Blackwell architectures, new GPUs (such as H100 and B200) now provide native support for FP8, which has emerged as a promising precision due to its potential for larger computational and memory efficiency.
@@
Table 5: Fine-tuning recipe. Our recommendation is to tune the LLM with either LoRA or QLoRA and to tune ViT’s layer normalization (LN) layers with a much smaller learning rate. This setup achieves competitive accuracy and is also the most memory- and compute-efficient. All experiments use a batch size of 1 with gradient checkpointing disabled, and throughput is measured on a single NVIDIA A100 80GB GPU. For settings with {1,5,10,50}, we select the learning rate ratio from this set that gives the best results for each benchmark. “FT-5” refers to the average accuracy across AITZ [36], ALFRED [37], nuScenes [38], PathVQA [39], and Widget Caption [40].
ViT LLM Memory (GB) Throughput (iter/s) $\texttt{LR}_{\texttt{LLM}}/\texttt{LR}_{\texttt{ViT}}$ Accuracy (FT-5)
LoRA LoRA 20.1 3.4 1 69.2
{1,5,10,50} 71.8
LN LoRA 19.2 4.5 1 63.5
{1,5,10,50} 71.4
FT LoRA 21.9 4.2 1 64.0
{1,5,10,50} 70.1
LoRA QLoRA 11.1 2.6 1 63.0
{1,5,10,50} 70.8
LN QLoRA 10.2 3.1 1 62.7
{1,5,10,50} 70.9
FT FT 63.5 6.1 1 77.7
@@ 2.3 Efficient Fine-Tuning
When fine-tuning the vision encoder (ViT) and language model (LLM) together using PEFT methods, we observe that the learning rate should be set differently for VE and LLM: the learning rate for the ViT part will be 5-50$\times$ smaller than that for the LLM part. On the other hand, we also observe that fine-tuning the vision encoder with Layernorm can achieve comparable performance as LoRA (Table. 5) while being more computationally efficient: it can reduce the training time by 25% compared to applying LoRA for the vision encoder. With the curated configuration setup, NVILA can be quickly fine-tuned to various downstream tasks under 24 GB memory with on-par performance.
Table 6: Quantization recipe. While W4A16 quantization on LLM backbone may introduce small accuracy drop, W8A8 quantization on ViT is nearly lossless.
ViT LLM AI2D MMMU VideoMME TTFT (s)
FP16 FP16 91.0 50.7 63.9 0.90
FP16 W4A16 90.9 49.2 62.0 0.77
W8A8 W4A16 90.9 49.3 62.1 0.65
@@ 2.4 Efficient Deployment
We develop a specialized inference engine with quantization techniques to efficiently deploy NVILA. The inference process is divided into two phases: prefilling and decoding. In the compute-bounded prefilling stage, we first apply token compression techniques (Section 2.1) to reduce the inference workload for LLM backbone, after which the vision tower becomes the primary bottleneck, accounting for over 90% of the prefilling latency. To tackle this, we implement W8A8 quantization for the vision tower to reduce NVILA’s Time-To-First-Token (TTFT) in this compute-bounded stage. For the memory-bounded decoding stage, we follow AWQ [41] for W4A16 quantization of the LLM backbone to accelerate. We further optimize the original AWQ implementation by introducing FP16 accumulation to the W4A16 GEMM kernels, resulting to a total 1.7$\times$ kernel speedup without compromising accuracy. We attach a detailed comparison in Figure. 5.
3 Experiments
3.1 Training Details
@@ 3.2 Accuracy Results
Table 8: Image benchmarks. We mark the best performance bold and the second-best underlined.
@@
test test test test testmini val test pro image val testdev
GPT-4o 94.2 85.7 92.8 79.2 63.8 69.1 64.7 51.9 75.4 76.2 77.4 78.7
Claude 3.5 Sonnet 94.7 90.8 85.2 74.3 67.7 68.3 63.7 51.5 60.1 74.1 70.7
Gemini 1.5 Pro 94.4 87.2 93.1 81.0 63.9 62.2 57.6 43.5 70.4 78.7 80.2
LLaVA-1.5 7B 55.5 17.8 28.1 25.8 25.6 35.7 54.8 66.1 58.2 78.5
VILA-1.5 8B 76.6 52.7 40.6 25.9 36.7 38.6 32.7 52.7 73.8 68.5 83.0
Cambrian-1 8B 73.0 73.3 77.8 41.6 49.0 42.7 64.2 74.7 71.7 81.2
Florence-VL 8B 74.2 74.7 84.9 51.7 55.5 43.7 64.2 74.9 74.2 84.7
LLaVA-OneVision 8B 81.4 80.0 87.5 68.8 63.2 48.8 42.8 24.1 66.3 75.4 78.3 84.0
Llama 3.2 11B 91.9 83.4 88.4 51.5 50.7 75.2
InternVL2 8B 83.8 83.3 91.6 74.8 58.3 51.2 42.6 29.0 64.2 76.2 77.4 76.7
Qwen2-VL 8B 83.0 83.0 94.5 76.5 58.2 54.1 46.6 30.5 70.1 76.0 84.3 82.9
NVILA-Lite 8B 91.0 84.8 91.7 67.9 64.5 50.7 45.7 26.5 65.6 76.3 78.1 85.0
NVILA 8B 92.3 86.1 93.7 70.7 65.4 49.9 44.4 27.8 68.6 76.5 80.1 85.4
LLaVA-1.5 13B 61.1 18.2 30.3 29.4 27.7 37.0 55.3 68.2 61.3 80.0
VILA-1.5 13B 79.9 59.5 58.6 30.4 42.7 37.9 33.6 57.5 72.6 65.0 82.8
Cambrian-1 13B 73.6 73.8 76.8 48.0 40.0 63.0 74.4 72.8
Pixtral 12B 79.0 81.8 90.7 50.8 58.0 52.5 65.4 75.7 80.2
NVILA-Lite 15B 92.0 81.8 90.6 69.3 61.7 58.7 51.8 33.7 67.1 75.6 77.3 83.7
NVILA 15B 94.1 86.9 94.0 73.5 66.1 56.7 51.8 33.8 69.5 76.6 80.0 84.8
LLaVA-NeXT 34B 46.5 48.1 44.5 22.9 75.9 69.5 83.7
Cambrian-1 34B 79.7 75.6 75.5 46.0 53.2 49.7 67.8 75.3 76.7 83.8
VILA-1.5 40B 88.9 67.8 58.6 38.4 49.3 51.9 46.9 25.0 60.8 69.1 73.6 84.3
InternVL2 40B 87.1 86.2 93.9 78.7 63.7 55.2 47.4 34.2 71.8 78.2 83.0
LLaVA-OneVision 72B 85.6 83.7 91.3 74.9 67.5 56.8 52.3 31.0 71.9 75.4 80.5 85.2
NVLM-D-1.0 78B 94.2 86.0 92.6 65.2 59.7 54.6 69.7 82.1 85.4
Llama 3.2 90B 92.3 85.5 90.1 57.3 60.3 39.5
@@ 3.2.1 Image Benchmarks
Our NVILA performs comparably to top open-source models in each size category, including Qwen2-VL [5], InternVL [3], and Pixtral. For general visual question answering tasks (ChartQA, DocVQA, InfoVQA, TextVQA, VQAv2, Seed), NVILA-8B and NVILA-15B achieve competitive or even better results compared to proprietary models (GPT-4o, Gemini). In science-related benchmarks (AI2D), NVILA-8B achieves state-of-the-art performance among open-source models. When scaling to 15B, NVILA demonstrates competitive performance with proprietary models.
Furthermore, on reasoning and knowledge benchmarks such as MMMU, RealworldQA, and MathVista, scores improve more when the model size increases. For benchmarks that require OCR capability such as TextVQA, AI2D, ChartQA, DocVQA, InfoVQA, 8B model can also do a great job. We also show a few qualitative examples in Figure. LABEL:fig:qualitative_example to demonstrate the OCR, reasoning, and multi-image capability of the NVILA model.
@@ 3.2.2 Video Benchmarks
We evaluate our models on a range of video understanding benchmarks [57, 58, 59, 60], spanning short videos of a few seconds to longer videos up to an hour in duration. Table 9 presents the performance of NVILA compared to baseline models [61, 62, 5, 4, 63, 19]. NVILA features long-context capability and can process up to 256 frames. With the scale-then-compress design, NVILA-8B achieves impressive results, setting new state-of-the-art performance across all benchmarks. Notably NVILA reaches performance levels comparable to GPT-4o mini with only 8B parameters and outperforms many larger models.
Table 9: Video benchmarks.
@@
InternVL2 8B 64 54.6 64.0 65.8 56.3 59.3
LLaVA-OneVision 8B 32 56.6 56.5 64.7 56.7 79.4 58.2 61.5
Oryx-1.5 8B 128 56.3 67.5 67.6 81.8 58.8 64.2